You are in the 35% tax bracket. The apartment rents for $1800 per month. Your monthly mortgage payments would be ​2100​, of which an average of ​$1900 per month goes toward interest during the first year. Determine whether renting or buying is cheaper in terms of monthly payments during the first year. Assume you are itemizing deductions.
Answer by Boreal(15235)

Renting costs are 1800*12=$21,600
mortgage payments are 2100*12=$25,200
1900*12*0.35=deduction=$7980
the outlay with a mortgage is $25200-$7980=$17,220
buying and mortgage would be less expensive.